# Software Performance Engineering for Foundation Model-Powered Software (FMware)

### v1

@@ III-C Challenge 3: The complexity of performance tuning and optimization of FMware
For controlling the self-exploration behaviour, one approach is to first record the self-exploration and replay to reproduce the performance issue. Chen et al. [68] proposed to reproduce the model training process with a record and replay mechanism. Similar ideas can be applied to FMware inference, e.g., the decoding process for executing each FM invocation and other traditional software executions can be recorded and replayed as an agent explores, facilitating the analysis and debugging of performance issues in a reproducible manner.
III-D Challenge 4: The complexity of deploying FMware
@@
Several key decisions must be carefully considered when deploying FMware into production, as it must meet service-level agreements (SLAs). While many studies focus on ensuring the SLAs at the model level [69, 70, 71] , these efforts alone are insufficient to guarantee the FMware meets application-level SLAs, since models are only part of the entire software system. Driven by such requirements, we identify challenges in three dimensions described as follows:
@@
Selecting optimal deployment options when hosting FMware: Unlike deploying traditional software, FMware deployment requires higher computation costs due to the invocation of FMs, which often involve the usage of specialized accelerators. There are three types of deployment options for FMware: API-based deployment, rented cloud instances, and on-premise self-hosting. In certain cases, these deployment options can also be jointly leveraged. API-based deployment follows a pay-as-you-go mechanism. Examples are OpenAI-compatible APIs by proprietary model providers [72] or Anyscale Model Endpoints API [73]. Developers send HTTP requests to served models to retrieve the generated tokens. While it is the simplest way to set up, the performance of API-based deployments solely relies on the API provider and can sometimes be unpredictable or unreliable [74]. Rented cloud instances refer to renting computation resources from cloud service providers. Developers could either rent compute instances for dedicated purposes (e.g., AWS EC2 G5 [75]) or in a serverless way [76]. This option provides flexibility and can absorb spikes in request volume, as these platforms usually provide autoscaling mechanisms. However, it requires DevOps engineers to configure based on the computation requests. A common challenge is low hardware utilization resulting in unnecessary costs. On-premise hosting means that a person or an organization procures physical or managed private cloud clusters which are dedicated to them. Such an option provides the maximum flexibility and control over hardware. At the same time, extensive engineering efforts are needed to guarantee the optimal usage of these hardware to satisfy the needs for multi-tenancy, as the resources usually need to be shared to cover the costs. The three above-mentioned deployment options have pros and cons. Hence, developers need to balance the degree of control over hardware, the costs, the utilization, and the expected application performance.
@@
For API-based deployment, existing practices attempted mixed use of small and large FMs for latency reduction. Both BiLD [77] and Minions [78] showed that smaller models are effective in latency reduction (dropping to 50% in the case of BiLD) with little-to-no output generation quality compromise. To control the unpredictability of APIs, Wang et al. [79] studied the request and response token length distributions of ChatGPT and GPT4 models at the API level, and proposed that this trace can be used for optimizing serving systems to become “workload-aware”.
@@
Rented cloud and on-premise hosting require a balance of performance and cost. Griggs et al. [80] remarked that the optimal GPU for cost-efficiency in running an FM varies and largely depends on the size of requests being processed. They showed that up to 77% cost reduction is possible with heterogeneous GPU type selection at the time of deployment. Several industrial solutions like run:AI [81] and Apache YuniKorn [82] improve the utilization of accelerators through advanced scheduling features. However, none of the current solutions consider application-level SLA requirements. Future work might explore the possibility of hosting FMware in a hybrid way. In combination with application-level SLA-aware scheduling [83] and scaling algorithms [80, 84] , more research is needed to achieve the optimal performance and lowest costs for deploying FMware.
@@
Even within one process such as inference, space separation at this finer granularity is shown to be effective. Disaggregation was applied by Hu et al. [85] to separate prefill and decode instances, improving the performance over cost metric by 2.4x. Memory capacity is also a scarce resource in this scenario, and the variability of prompt and output tokens for one inference process leads to variability in leftover memory for other processes. This is largely due to the KV cache memory used for each output token, further complicating memory allocation and scheduling strategies to attain maximum throughput in serving systems. Cheng et al. [86] proposed a “Wasted Memory Access” (WMA) metric to accurately predict memory consumption, so that corresponding memory required at certain batch sizes of inference requests can be leveraged for smart scheduling decisions.
When dealing with complex cognitive architectures, these issues are further amplified. Unlike traditional runtime systems where resources need to be predefined, Agentware requires dynamic allocation, as agents operate autonomously without defined code paths and share runtime resources on the fly. Under a resource-constrained environment, this can lead to contention and interference among agents, impacting other processes’ performance. New “OS-like” architectures have been proposed, where FMs act as the kernel to govern access to shared hardware resources and services [87, 27], to make agent completion latency predictable. In addition, Mei et al. [87] show that isolation between modules in the LLM Kernel is the key to preventing resource conflicts with the rest of the system, and ensuring optimal access to resources and services when agents execute tasks.
@@
Deploying multi-tenant FMware efficiently: Current infrastructures and hardware accelerators are too expensive to dedicate to a single FMware, necessitating a need for multi-tenant optimization objectives. It is often economical to have a cluster shared among multiple deployments, which introduces the challenge of multi-tenancy. The primary objective is to maximize the cluster-level hardware utilization and efficiency of the shared hardware across all FMware deployments while trying to meet every tenant’s performance goals; each tenant may have a different volume of users and SLA requirements. In some scenarios, there can be conflicting performance requirements when diverse types of FMware are co-located, as such a uniform cluster-wide policy is not optimal. Lazuka et al. [88] illustrated the difficulty of simultaneously meeting a low-latency SLA required by chatbot workloads and a high throughput SLA required for text summarization workloads, within the same serving system.
@@
One way to share the cluster is through sharing served FMs across different FMware and keeping models persisted in memory to avoid model loading costs, as well as reducing the occupied memory. Selecting batch sizes of requests to optimize for both latency and throughput across many deployed FMs becomes challenging. This is particularly important for Promptware, where multiple FMs may be used in a pipeline or workflow of invocations, under end-to-end SLA constraints. Tan et al. [83] showed that topology-aware batching at the application level can achieve a latency reduction of up to 19% under multi-query workload scenarios. The batching is performed based on workflow dependency of multiple requests to meet multiple SLAs simultaneously, and a batch size is selected to be most efficiently processed by the execution engine. However, scaling strategies when existing resources are not enough to meet request SLAs were not considered in the scope of this work.
@@
In addition to batching, optimally routing requests to compute resources is another critical aspect. Lin et al. [89] demonstrated that routing requests with a common prefix to the same inference instance maximizes the reuse of existing KV cache, thus achieving high locality and utilization. Their proposed Semantic Variable is a declarative approach at the FMware level for schedulers to be optimized around user intent. These strategies also help improve memory usage predictability and overall performance stability, since the size of KV cache used is known during the routing process, instead of allocating the cache anew. To minimize data movement, Sun et al. [19] proposed a live migration mechanism to enable runtime rescheduling, resulting in 26x lower latency at the prefill phase. This is accomplished using a two-level global and instance-level scheduler. The former coordinates request dispatching, migrations, and autoscaling actions, and the latter reports the memory load and virtual usage of each instance back to the global scheduler. Both works are limited however as they do not simultaneously consider SLA aspects for scaling decisions at the FMware level, only at each inference request from a “Model-as-a-Service” understanding.
@@
Existing studies remain siloed, lacking integration among model-level, multi-tenant optimization and cluster elasticity. Future research should explore advanced methods for understanding user behaviour, intent [89], and interference sensitivity. Expanding beyond “cluster-in-a-vacuum”-scale to internet-scale awareness [74] would enable intelligent, flexible policies to anticipate performance degradation for each tenant. FMware serving should jointly consider scheduling (assigning requests to model replicas) and resource allocation (determining the number of replicas to deploy) to meet multiple performance objectives at the application level. In the next section, we discuss our system design to achieve these goals.
IV Our vision towards an SLA-aware FMware Runtime
@@
In this section, we present our vision and reference architecture for a performance-oriented runtime designed to serve multi-tenant FMware, specifically Promptware. Our approach prioritizes SLAs as the central design principle, aiming to satisfy each FMware’s SLA requirements while optimizing cluster-level hardware utilization.
@@
Our architecture is tailored for Promptware, where the application is represented as DAGs. The nodes of the graph represent tasks and the edge represents the control flow dependencies (e.g., sequential or conditional branching). To simplify the description, we treat each task as an invocation to an FM. However, in practice, the tasks can also take other forms, such as regular code execution or external API calls served outside our runtime. The simplified architecture and the core components are illustrated in Figure 2.
@@
Figure 2: The simplified architecture and components of FMware Runtime
Our runtime can be deployed at an on-premise cluster or a rented cloud cluster as long as we have full access to the machines. Each machine is equipped with the same accelerator setup (e.g., 8 Ascend 910B NPUs [90] per machine). For each FMware workflow that is going to be deployed at FMware Runtime, we will decompose the DAG representation into relevant FM invocations. For example, one workflow request could be represented as a sequential invocation of Model A, B, and C. Note that the model can be reused in multiple FMware as we have discussed in Section III-D. To meet performance requirements, we present the four core components of our reference architecture:
@@
  • Profiler. This component is responsible for calculating the estimated latency for the invocation of each type of FM as well as the memory consumption. We control a default number of tokens to eliminate the non-deterministic behaviour during profiling. The referenced latency can be used as an estimation of how much time each task should consume so that it will not impact the end-to-end latency goal. We refer to this concept as “slack” - a portion of the total SLA allocated to each task. Profiling is done offline, and each FM is profiled only once.
@@
  • Resource Provisioner. This component makes decisions about allocating or releasing accelerator resources based on the risk of SLA violations. If SLAs are at risk, it checks for available resources to spin up new FM replicas. Conversely, if a replica remains idle, it will release the unused resources.
  • Replica Router. This component is responsible for routing the model requests into the model replicas. As the runtime receives a larger load of requests, more replicas will be created by the Resource Provisioner to ensure SLAs. It is crucial to have an intelligent router to decide which replica should serve an incoming request, based on how much “slack” it still has. For example, if earlier tasks consume less time than expected, requests can be routed to replicas with longer queues to balance the load. If the upstream model requests in an application take time that is shorter than expected, the incoming model request can then be routed to a replica that has more requests queueing, as it could wait a bit longer and vice versa. The Resource Provisioner will work with this component to monitor SLA compliance and make joint decisions together.
@@
  • Cluster. This component handles the execution of commands from the Resource Provisioner and Replica Router. It manages cross-node communication and data movement such as loading model weights.
We have developed a prototype system based on the design described above, which has already been deployed internally in production in a cloud environment. Initial evaluation results demonstrate that our SLA-aware FMware Runtime outperforms established open source solutions like Ray Serve [91], which can also handle routing requests and scaling up instances when needed. However, without considering SLA constraints, existing solutions would experience a higher SLA violation rate as the request load increases. In the future, We plan to extend the system to cope with other aforementioned challenges.
V Conclusion

### v2

@@ 3.4. Challenge 3: The complexity of performance tuning and optimization of FMware
For controlling the self-exploration behaviour, one approach is to first record the self-exploration and replay to reproduce the performance issue. Chen et al. (Chen et al., 2022) proposed to reproduce the model training process with a record and replay mechanism. Similar ideas can be applied to FMware inference, e.g., the decoding process for executing each FM invocation and other traditional software executions can be recorded and replayed as an agent explores, facilitating the analysis and debugging of performance issues in a reproducible manner.
3.5. Challenge 4: The complexity of deploying FMware
@@
Several key decisions must be carefully considered when deploying FMware into production, as it must meet service-level agreements (SLAs). While many studies focus on ensuring the SLAs at the model level (Oh et al., 2024; Wu et al., 2023; Gujarati et al., 2020) , these efforts alone are insufficient to guarantee the FMware meets application-level SLAs, since models are only part of the entire software system. Driven by such requirements, we identify challenges in three dimensions described as follows:
@@
Selecting optimal deployment options when hosting FMware: Unlike deploying traditional software, FMware deployment requires higher computation costs due to the invocation of FMs, which often involve the usage of specialized accelerators. There are three types of deployment options for FMware: API-based deployment, rented cloud instances, and on-premise self-hosting. In certain cases, these deployment options can also be jointly leveraged. API-based deployment follows a pay-as-you-go mechanism. Examples are OpenAI-compatible APIs by proprietary model providers (Google, ) or Anyscale Model Endpoints API (Anyscale, 2024). Developers send HTTP requests to served models to retrieve the generated tokens. While it is the simplest way to set up, the performance of API-based deployments solely relies on the API provider and can sometimes be unpredictable or unreliable (Perez, 2024). Rented cloud instances refer to renting computation resources from cloud service providers. Developers could either rent compute instances for dedicated purposes (e.g., AWS EC2 G5 (Amazon, )) or in a serverless way (Lu, ). This option provides flexibility and can absorb spikes in request volume, as these platforms usually provide autoscaling mechanisms. However, it requires DevOps engineers to configure based on the computation requests. A common challenge is low hardware utilization resulting in unnecessary costs. On-premise hosting means that a person or an organization procures physical or managed private cloud clusters which are dedicated to them. Such an option provides the maximum flexibility and control over hardware. At the same time, extensive engineering efforts are needed to guarantee the optimal usage of these hardware to satisfy the needs for multi-tenancy, as the resources usually need to be shared to cover the costs. In terms of cost, API-based deployment can incur generation costs of roughly \$35k per month or \$420k per year, making it the largest single cost component for organizations (Roux, 2023). Cloud-based LLM deployments provide flexible pay-as-you-go pricing, but costs can become unpredictable at high usage levels and may reach 2-3 times the cost of on-premise solutions for large-scale operations (Latitude, 2025). On-premise self-hosting involves significant upfront investment but delivers more predictable costs and can yield 30-50% savings over three years when hardware utilization consistently exceeds 60-70% (Latitude, 2025). The three above-mentioned deployment options have pros and cons. Hence, developers need to balance the degree of control over hardware, the costs, the utilization, and the expected application performance.
@@
For API-based deployment, existing practices attempted mixed use of small and large FMs for latency reduction. Both BiLD (Kim et al., 2024) and Minions (Wang et al., 2024b) showed that smaller models are effective in latency reduction (dropping to 50% in the case of BiLD) with little-to-no output generation quality compromise. To control the unpredictability of APIs, Wang et al. (Wang et al., 2024d) studied the request and response token length distributions of ChatGPT and GPT4 models at the API level, and proposed that this trace can be used for optimizing serving systems to become “workload-aware”.
@@
Rented cloud and on-premise hosting require a balance of performance and cost. Griggs et al. (Griggs et al., 2024) remarked that the optimal GPU for cost-efficiency in running an FM varies and largely depends on the size of requests being processed. They showed that up to 77% cost reduction is possible with heterogeneous GPU type selection at the time of deployment. Several industrial solutions like run:AI (Sharon, 2024) and Apache YuniKorn (YuniKorn, 2024) improve the utilization of accelerators through advanced scheduling features. However, none of the current solutions consider application-level SLA requirements. Future work might explore the possibility of hosting FMware in a hybrid way. In combination with application-level SLA-aware scheduling (Tan et al., 2024) and scaling algorithms (Griggs et al., 2024; Anyscale, ) , more research is needed to achieve the optimal performance and lowest costs for deploying FMware.
@@
Even within one process such as inference, space separation at this finer granularity is shown to be effective. Disaggregation was applied by Hu et al. (Hu et al., 2024) to separate prefill and decode instances, improving the performance over cost metric by 2.4x. Memory capacity is also a scarce resource in this scenario, and the variability of prompt and output tokens for one inference process leads to variability in leftover memory for other processes. This is largely due to the KV cache memory used for each output token, further complicating memory allocation and scheduling strategies to attain maximum throughput in serving systems. Cheng et al. (Cheng et al., 2024) proposed a “Wasted Memory Access” (WMA) metric to accurately predict memory consumption, so that corresponding memory required at certain batch sizes of inference requests can be leveraged for smart scheduling decisions.
When dealing with complex cognitive architectures, these issues are further amplified. Unlike traditional runtime systems where resources need to be predefined, Agentware requires dynamic allocation, as agents operate autonomously without defined code paths and share runtime resources on the fly. Under a resource-constrained environment, this can lead to contention and interference among agents, impacting other processes’ performance. New “OS-like” architectures have been proposed, where FMs act as the kernel to govern access to shared hardware resources and services (Mei et al., 2024; Li et al., 2024), to make agent completion latency predictable. In addition, Mei et al. (Mei et al., 2024) show that isolation between modules in the LLM Kernel is the key to preventing resource conflicts with the rest of the system, and ensuring optimal access to resources and services when agents execute tasks.
@@
Deploying multi-tenant FMware efficiently: Current infrastructures and hardware accelerators are too expensive to dedicate to a single FMware, necessitating a need for multi-tenant optimization objectives. It is often economical to have a cluster shared among multiple deployments, which introduces the challenge of multi-tenancy. The primary objective is to maximize the cluster-level hardware utilization and efficiency of the shared hardware across all FMware deployments while trying to meet every tenant’s performance goals; each tenant may have a different volume of users and SLA requirements. In some scenarios, there can be conflicting performance requirements when diverse types of FMware are co-located, as such a uniform cluster-wide policy is not optimal. Lazuka et al. (Łazuka et al., 2024) illustrated the difficulty of simultaneously meeting a low-latency SLA required by chatbot workloads and a high throughput SLA required for text summarization workloads, within the same serving system.
@@
One way to share the cluster is through sharing served FMs across different FMware and keeping models persisted in memory to avoid model loading costs, as well as reducing the occupied memory. Selecting batch sizes of requests to optimize for both latency and throughput across many deployed FMs becomes challenging. This is particularly important for Promptware, where multiple FMs may be used in a pipeline or workflow of invocations, under end-to-end SLA constraints. Tan et al. (Tan et al., 2024) showed that topology-aware batching at the application level can achieve a latency reduction of up to 19% under multi-query workload scenarios. The batching is performed based on workflow dependency of multiple requests to meet multiple SLAs simultaneously, and a batch size is selected to be most efficiently processed by the execution engine. However, scaling strategies when existing resources are not enough to meet request SLAs were not considered in the scope of this work.
@@
In addition to batching, optimally routing requests to compute resources is another critical aspect. Lin et al. (Lin et al., 2024) demonstrated that routing requests with a common prefix to the same inference instance maximizes the reuse of existing KV cache, thus achieving high locality and utilization. Their proposed Semantic Variable is a declarative approach at the FMware level for schedulers to be optimized around user intent. These strategies also help improve memory usage predictability and overall performance stability, since the size of KV cache used is known during the routing process, instead of allocating the cache anew. To minimize data movement, Sun et al. (Sun et al., 2024a) proposed a live migration mechanism to enable runtime rescheduling, resulting in 26x lower latency at the prefill phase. This is accomplished using a two-level global and instance-level scheduler. The former coordinates request dispatching, migrations, and autoscaling actions, and the latter reports the memory load and virtual usage of each instance back to the global scheduler. Both works are limited however as they do not simultaneously consider SLA aspects for scaling decisions at the FMware level, only at each inference request from a “Model-as-a-Service” understanding.
@@
Existing studies remain siloed, lacking integration among model-level, multi-tenant optimization and cluster elasticity. Future research should explore advanced methods for understanding user behaviour, intent (Lin et al., 2024), and interference sensitivity. Expanding beyond “cluster-in-a-vacuum”-scale to internet-scale awareness (Perez, 2024) would enable intelligent, flexible policies to anticipate performance degradation for each tenant. FMware serving should jointly consider scheduling (assigning requests to model replicas) and resource allocation (determining the number of replicas to deploy) to meet multiple performance objectives at the application level. In the next section, we discuss our system design to achieve these goals.
4. Our vision towards an SLA-aware FMware Runtime
4.1. FMware Runtime Design
@@
In this section, we present the design, architecture, and implementation details of our performance-oriented runtime (FMware Runtime), which is built to serve multi-tenant FMware. Our approach prioritizes SLAs as the central design principle, aiming to satisfy each FMware’s SLA requirements while optimizing cluster-level hardware utilization.
@@
Our design is specifically tailored for Promptware, not Agentware. Agentware’s dynamic, autonomous control flow and evolving behavior introduce complexity (such as non-deterministic execution and recursive self-correction) that falls outside the scope of this study. Instead, our focus is on Promptware, where the application is represented as DAGs. The nodes of the graph represent tasks and the edge represents the control flow dependencies (e.g., sequential or conditional branching). To simplify the description, we treat each task as an invocation to an FM. However, in practice, the tasks can also take other forms, such as regular code execution or external API calls served outside our runtime.
Our runtime can be deployed on an on-premise cluster or a rented cloud cluster, provided we have full access to the machines. While individual machines are equipped with the same accelerator setup (e.g., 8 Ascend 910B NPUs (Liang, 2020) per machine), the runtime is fully aware of and leverage the cluster-wide heterogeneity. For each FMware workflow deployed on FMware Runtime, the workflow is first represented as a DAG. During execution, this DAG is decomposed into concrete model requests, meaning each node that involves an FM invocation is translated into one or more calls to the relevant model(s). For example, a workflow request could be represented as a sequential DAG with nodes invoking Model A → Model B → Model C, where the outputs of one node can serve as inputs to downstream nodes. Models can also be reused across multiple workflows, as discussed in Section 3.5, allowing shared FM resources to serve different workflow requests efficiently.
Figure 3. The simplified architecture and components of FMware Runtime
@@
The simplified architecture and the core components of FMware Runtime are illustrated in Figure 3. To meet performance requirements, we present the four core components of our system architecture:
@@
  • Profiler. This component is responsible for offline profiling of each FM on each available hardware type. It calculates the estimated inference latency (given a default number of tokens to ensure determinism) as well as memory consumption for each type of FM invocation. This profile data is crucial, as it provides referenced measurements used by the runtime. For example, the referenced latency can be used as an estimation of how much time each task should consume so that it will not impact the end-to-end latency goal. It also translates the end-to-end workflow SLO into per-task time budgets, which we refer to as “slack” – a portion of the total SLO allocated to each task. Profiling is done offline, and each FM is profiled only once.
@@
  • Resource Provisioner. This component acts as the coarse-grained control for SLO adherence. It makes strategic decisions about allocating or releasing accelerator resources based on the monitored risk of SLA violations. If workflows are frequently exhausting their slack and risk missing end-to-end SLOs, the Resource Provisioner checks for available cluster resources to spin up new FM replicas. Conversely, if a replica remains idle for a configured period, it will be terminated to free up resources.
  • Replica Router. This component is the fine-grained control and provides the core decision-making function of FMware Runtime. It is responsible for routing each individual model request to a specific model replica. As the Resource Provisioner creates more replicas (potentially on different hardware types), the Replica Router must intelligently decide which replica should serve an incoming request. To achieve this, the Replica Router implements a novel, SLO-aware scheduling algorithm (Algorithm 1). Unlike static routing (e.g., round-robin), our scheduling mechanism dynamically considers user SLOs (via slack) and the heterogeneous card types available. The algorithm’s core is an objective latency cost function that decides the optimal hardware configuration (i.e., which specific replica on which card type) to use for each node invocation in the workflow. This cost function estimates the total latency for a request on a specific hardware type by combining three key factors: (a) Estimated Inference Latency $lx$: The base execution time for the model on that specific hardware, derived from the Profiler; (b) Queuing Delay $lq$: The estimated time the request would spend waiting in the target replica’s queue; (c) In-flight Request Delay $lr$: The processing time remaining for requests already being executed (in-flight) by that replica. The final estimated latency $l_{est}$ is the sum of these factors: $l_{est}=lx+lq+lr$. For every incoming request, the Replica Router iterates through all possible configurations (i.e., all available replicas across all hardware types), calculates $l_{est}$, and selects the configuration that minimizes this estimated latency. This ensures that each request is routed to the replica most likely to meet its slack requirement, thereby protecting the end-to-end workflow SLO.
@@
  • Cluster. This component is the execution layer that handles commands from the Resource Provisioner and Replica Router. It manages cross-node communication, data movement, and the lifecycle of model replicas (e.g., loading model weights into accelerator memory).
Algorithm 1 Scheduling Algorithm
1:$n$: node to be scheduled
2:$C$: list of profiled configurations for the node
3:slackslack: estimated slack for the node
4:procedure FWOSchedule($n,C$, $s$)
5:  costs_targetcosts\_target = { }, costs_greedycosts\_greedy = { }
6:  for $c1$ in $C$ do
7:   lxlx = ProfiledLatency($c1$)
8:   cardcard = c1.card_typec1.card\_type
9:   lqlq = QueingDelay(cardcard)
10:   lrlr = InflightRequestsDelay(cardcard)
11:   costcost = $lx+lq+lr$
12:   if $cost<slack$ then
13:     costs_target.put(cost)costs\_target.put(cost)
14:   else
15:     costs_greedy.put(cost)costs\_greedy.put(cost)
16:   end if
17:  end for
18:  configconfig = Min(costs_target,costs_greedycosts\_target,costs\_greedy)
19:  n.set_config(confign.set\_config(config)
20:end procedure
The design of FMware Runtime prioritizes robust end-to-end SLO enforcement, leading to several key trade-offs. First, we applied a dynamic, SLO-aware scheduling algorithm over simpler static load balancers (like Ray Serve’s baseline). While static approaches are simpler to implement, they are SLO-agnostic and poorly handle hardware heterogeneity, leading to SLO violations. We traded this simplicity for a higher scheduling complexity in our Replica Router, which is essential for fine-grained control. Secondly, our Replica Router operates within a logically centralized control plane to maintain a global view of all workflows. The trade-off is that a decentralized design might offer greater scheduler scalability, but it sacrifices the global context necessary to make optimal decisions for end-to-end SLOs. Finally, within our Replica Router, we chose a cost function considering three aspects ($l_{est}=lx+lq+lr$) instead of simpler heuristics such as the shortest queue. While computationally cheaper, heuristics ignore crucial factors like hardware differences (e.g., a short queue on a slow card) and in-flight request state. We traded this lower computational cost for the superior accuracy of our approach, which provides a holistic latency estimate critical for making correct scheduling decisions in a heterogeneous environment.
Discussion. The FMware Runtime architecture is explicitly designed to solve the complexities of Challenge 3 (Performance Tuning and Optimization) and Challenge 4 (Deployment). The Profiler addresses Challenge 3 by converting the complex, multi-knob optimization problem into a singular slack metric, simplifying performance optimization. It also provides the foundational profiling data needed to address the evolving nature of performance targets. The Resource Provisioner tackles Challenge 4’s multi-process multi-tenancy issue by providing dynamic elasticity, ensuring balanced resource usage by scaling resources based on risk of SLA violation. Finally, the Replica Router provides the core solution: its SLO-aware scheduling algorithm leverages the cost function to algorithmically optimize performance (Challenge 3) and make efficient, real-time routing decisions required for stable multi-tenant deployment (Challenge 4) across heterogeneous hardware.
4.2. Quantitative Evaluation
To validate our proposed design of FMware Runtime, we have developed and deployed a complete implementation of the FMware Runtime internally in a production cloud environment. We conduct a quantitative evaluation for the SLA-aware FMware Runtime, comparing its performance against Ray Serve (Ray Team, 2024), a widely-adopted open source solution for FM serving that can also handle routing requests and scaling up instances when needed. In a recent survey, the only multi-model serving framework specifically mentioned is Ray, which supports serving multiple LLMs, underscoring the limited set of directly comparable baselines (Xu et al., 2024).
4.2.1. Evaluation Setup
The evaluation was conducted on a heterogeneous computing cluster to realistically assess FMware Runtime’s capability to manage hardware resources under SLA. The cluster comprised two nodes with distinct accelerator card types: one node featured a faster card type, and the other a slower card type.
We deployed two distinct FMApp workflows to comprehensively test the system:
  1. 1.
    Retrieval-Augmented Generation (RAG) Application: This simple workflow involves an initial model performing information fetching and lightweight computation, followed by a second model executing a more complex, computation-heavy generation task.
  2. 2.
    Industry-based FMApp: This represents a customer-facing shopping chatbot workflow. It is characterized by conditional and model nodes, executing different DAG paths based on user input, potentially invoking up to five model calls.
For both workflows, the SLA target was estimated based on the Time To First Token (TTFT). The target was defined as the TTFT measured under optimal conditions (i.e., without hardware competition) plus an acceptable buffer time, setting a realistic user expectation for initial output delivery.
To ensure reproducible testing for controlled and deterministic workloads, we used a very low temperature setting and no additional parameters during the model’s generation, mitigating variations from random token generation. Conversely, for testing under realistic and nondeterministic workloads, we utilized a more typical temperature and realistic model parameter settings to accurately simulate real-world user experiences. Load testing was performed by sending requests following a Poisson distribution, simulating natural, variable user arrival rates.
4.2.2. Evaluation Result
We evaluated the performance of FMware Runtime against the Ray Serve baseline, focusing first on the latency aspect. The results demonstrate that FMware Runtime’s dynamic, profile-aware dynamic scheduling provides a significant reduction in latency under both deterministic and nondeterministic conditions. We evaluated two primary scheduling modes: a greedy mode, which prioritizes immediate resource optimization, and a target mode, which balances load to meet SLA targets.
Latency. Under controlled, deterministic workloads, FMware Runtime achieves a substantial reduction in the request execution latency before the system reaches saturation. Compared to the baseline, the greedy and target modes reduce the average request execution latency by 33.81% and 30.7%, respectively. This improvement stems from FMware Runtime’s performance-oriented scheduling. We analyzed the scheduling decisions in a 15-request test case. The baseline system, Ray Serve, routed only 30% of the model executions to the faster accelerator card. In contrast, FMware Runtime’s greedy mode assigned 70% of executions to the faster card, while the target mode assigned 36.7% to balance throughput. Even though Ray Serve provides features such as static fractional GPU and auto-scaling in a fixed range to ensure the cluster’s performance efficiency, the nondeterministic nature of FMApp execution makes it difficult to adopt a static scheduling mechanism. On the other hand, our profile-aware dynamic scheduling in FMware Runtime: (1) leverages the profiled performance metadata of model execution on different accelerator card types, and (2) uses a dynamic scheduling algorithm together with a real-time feedback loop mechanism to efficiently execute FMApp requests.
We also measured TTFT, a critical metric for user-perceived latency. FMware Runtime reduced the average TTFT by 33% (greedy mode) and 30% (target node) across workloads from 1 to 30 requests.
To simulate a more realistic environment, we introduced nondeterminism in model generation, where inference times varied even with an identical max new token setting. In this scenario, FMware Runtime’s robustness became evident. Table 1 shows FMware Runtime consistently outperformed the baseline, achieving a lower median TTFT across all tested request loads, from a single request up to system saturation. This result confirms that our profile-driven dynamic scheduling approach adapts effectively to the inherent variability of model inference in real-world conditions.
Table 1. Compare the median TTFT and goodput rate across different numbers of requests among Ray Serve, FMware Runtime in target and greedy modes.
System Median TTFT Median Goodput Rate
Ray Serve 153.6 0.53
FMware Runtime (greedy) 147.9 0.54
FMware Runtime (target) 126.0 0.61
Goodput. Besides latency reduction, FMware Runtime demonstrates better efficiency in terms of goodput rate, defined as the proportion of requests successfully completed within the specified SLA target — a metric to quantify effective request completion under SLA constraints (Wang et al., 2024e). In the same experiment, we achieved a significantly higher goodput rate compared to the baseline, with an average increase of 26.39% in the greedy mode and 15.27% in the target mode when serving our studied FMApp case before system saturation. This enhanced performance translates directly to increased capacity for SLA-compliant service. For instance, before engaging any scaling strategies, using a 70% goodput rate threshold as an acceptable service quality boundary, Ray Serve was capable of processing 7 requests. In contrast, FMware Runtime could process 8 requests in the target mode and 10 requests in the greedy mode. This indicates that FMware Runtime more efficiently utilizes hardware resources, achieving a better goodput rate through its dynamic resource management. Furthermore, Table 1 shows that FMware Runtime maintains its high goodput rate even in the aforementioned realistic, nondeterministic scenario.
5. Conclusion